the average velocity of CO2 at T1 Kelvin and maximum most probable velocity at T2 Kelvin is 9×10^4.

Grade:11

2 Answers

Ravleen Kaur
4 years ago
The question is incomplete. Please post one more time.
Vikas TU
4 years ago
Dear student 
Th question is incomplete , but if you want to find the alue of T1 and T2 . here is the solution 
Average velocity = √8RT/πM
and Most probable velocity = √2RT/M
Given –For CO2
Average velocity at T1 Most probable velocity at T2
= 9 * 104cm/sec = 9 * 104/100 m/sec.
= 9 * 102= m/sec.
∴ 9 * 102= √8 * 8.314 9 T1/3.14 *44 * 10-3 …..(A)
[Average velocity at T1K]
and 9 * 102= √2 * 8.314 * T2/44 * 10-3 …..(B)
[Most probable velocity at T2]
On solving, T1 = 1682.5 K, T2 = 2143.4 K